XBoard

Tim Mann Jochen Wiedmann

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991 Digital Equipment Corporation, Maynard, Massachusetts. All Rights Reserved Permission to use, copy, modify, and distribute this software and its documentation for any purpose and without fee is hereby granted, provided that the above copyright notice appear in all copies and that both that copyright notice and this permission notice appear in supporting documentation, and that the name of Digital not be used in advertising or publicity pertaining to distribution of the software without specific, written prior permission. Digital disclaims all warranties with regard to this software, including all implied warranties of merchantability and fitness. In no event shall Digital be liable for any special, indirect or consequential damages or any damages whatsoever resulting from loss of use, data or profits, whether in an action of contract, negligence or other tortious action, arising out of or in connection with the use or performance of this software. Enhancements copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 Free Software Foundation, Inc. Published by the Free Software Foundation 59 Temple Place - Suite 330 Boston, MA 02111-1307 USA Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU General Public License,” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the section entitled “GNU General Public License,” and this permission notice, may be included in translations approved by the Free Software Foundation instead of in the original English.

Introduction

This manual documents version 3.5.0 of XBoard. XBoard is a graphical chessboard that can serve as a user interface to the GNU Chess and Crafty chess programs, the Internet Chess Servers, electronic mail correspondence chess, or your own collection of saved games.

When this manual refers to "GNU Chess Mode", the information relates to both the GNU Chess and Crafty chess programs.

1 Other programs you can use with XBoard

Here are some other programs you can use with XBoard

1.1 GNU Chess

The GNU Chess program is available from: ftp://prep.ai.mit.edu/pub/gnu/gnuchess* (The version number changes fairly often, so I left it out. Just get whatever the latest version

is.)

You can use XBoard to play a game against GNU Chess, or to interface GNU Chess to an ICS.

1.2 Crafty

Crafty is a chess program written by Bob Hyatt (hyatt@cis.uab.edu). You can use XBoard to play a game against Crafty, hook Crafty up to an ICS, or use Crafty to interactively analyze games and positions for you.

To use Crafty with XBoard, give the -fcp option like this:

       xboard -fcp "crafty xboard"   ... (other options) ...

Crafty is a rapidly evolving program, which is good for its development, but sometimes bad for backwards compatibility. To make life simpler for everyone, an "XBoard-supported" version of Crafty is being provided. This version is guaranteed to work with XBoard. For details, see:

       http://fly.hiwaay.net/~frankm/crafty.html

Or, if you only have FTP access, look in the directory:

       ftp://gw2.maf.mobile.al.us/frankm/crafty

The "supported" version of Crafty includes some important changes that make it work better with XBoard. You can of course always grab the latest official version from:

       ftp://ftp.cis.uab.edu/hyatt/crafty.tar.gz

But please be aware of the caveats described on the Web page before reporting "bugs" to the XBoard developers.

1.3 zic2xpm

The “zic2xpm” program is used to import chess sets from the ZIICS(*) program into XBoard. “zic2xpm” is part of the XBoard distribution. ZIICS is available from:

ftp://ftp.onenet.net/pub/chess/DOS/ziics131.exe To import ZIICS pieces, do this:

1. Unzip ziics131.exe into a directory:

            unzip -L ziics131.exe -d ~/ziics

2. Use zic2xpm to convert a set of pieces to XBoard format.

For example, let’s say you want to use the FRITZ4 set. These files are named “fritz4.*” in the ZIICS distribution.

               mkdir ~/fritz4
               cd ~/fritz4
               zic2xpm ~/ziics/fritz4.*

3. Now, either give XBoard the “-xpm” option when starting up, e.g.:

               xboard -xpm ~/fritz4    (... other options ...)

Or, add this line to your .Xdefaults file:

               xboard*xpmDirectory: ~/fritz4

(*) ZIICS is a separate Copyrighted work of Andy McFarland (Zek on ICS). The “ZIICS pieces” are copyrighted works of their respective creators. Files produced by “zic2xpm” are for PERSONAL USE ONLY and may NOT be redistributed without explicit permission from the original creator(s) of the pieces.

2 How to configure, compile and install XBoard

XBoard is normally distributed as source, in a file with a name like ‘xboard-*.tar.gz’. You need to unpack this file and compile the results with a C compiler.

First, decide where you want to install the xboard executable program and documentation. The default is to put them under the ‘/usr/local’ directory tree. If you want to put them somewhere else, such as under your home directory, change the value in the ‘--prefix’ option below.

Use the following commands:

     gzip -cd xboard-*.tar.gz | tar -xvf -
     cd xboard-*/
     configure --prefix=/usr/local
     make
     make install

If you have any problems with these steps, see the file ‘INSTALL’ in the XBoard distribution for detailed instructions on options you can give to configure, and see the file ‘FAQ’ for answers to frequently asked questions.

If everything works, you can delete the whole ‘xboard-*’ directory after ‘make install’ is done.

3 Command options supported by XBoard

XBoard always runs in one of 4 possible modes:

xboard [options] starts XBoard as a GNU chess interface. As an interface to GNU Chess, XBoard lets you play a game against the machine, set up arbitrary positions, force variations, or watch a game between two machines.

xboard -ics [options] lets XBoard act as an interface to the Internet Chess Server (ICS). You can play against other ICS users, observe games they are playing, or review games that have recently finished. See Chapter 5 [ICS], page 25 All the “wild” chess variants on ICS are supported, including bughouse.

xboard -ncp [options] lets you use XBoard as a simple chessboard to play through games. It will read and write game files and allow you to play through variations manually. You can use it to browse games off the net or review GNU Chess, ICS, and correspondence games you have saved. These features are available at all times, even if you do not have access to GNU Chess or the ICS. If you want to pipe games into XBoard, use the supplied shell script ‘pxboard’. For example, from the news reader ‘xrn’, find a message with one or more games in it, click the Save button, and type ‘|pxboard’ as the file name.

cmail will use XBoard as an interface to electronic mail correspondence chess. See Chapter 7 [CMail], page 29.

Most of the XBoard options have both a long name and a short name. To turn a boolean option on or off from the command line, either give its long name followed by the value true or false (‘-longOptionName true’), or give just the short name to turn the option on (‘-opt’), or the short name preceded by ‘x’ to turn the option off (‘-xopt’). For options that take strings or numbers as values, you can use the long or short option names interchangeably.

Each option corresponds to an X resource with the same name, so you can set options in your ‘.Xdefaults’ file if you like. For options that have two names, the longer one is the name of the corresponding X resource; the short name is not recognized in ‘.Xdefaults’. To turn a boolean option on or off in your ‘.Xdefaults’ file, give its long name followed by the value true or false:

XBoard*longOptionName: true

3.1 Controlling the chess program

These options apply to both the GNU Chess and Crafty programs.

-tc or -timeControl minutes[:seconds] Each player begins with his clock set to the timeControl period. Default: 5 minutes. The additional options movesPerSession and timeIncrement are mutually exclusive.

-mps or -movesPerSession moves When both players have made movesPerSession moves, a new timeControl period is added to both clocks. Default: 40 moves.

-inc or -timeIncrement seconds If this option is specified, movesPerSession is ignored. Instead, after each player’s move, timeIncrement seconds are added to his clock. Use ‘-inc 0’ if you want to require the entire game to be played in one timeControl period, with no increment. Default: -1, which specifies movesPerSession mode.

-clock or -clockMode Determines whether or not to display the chess clocks. If clockMode is ‘false’, the clocks are not shown, but the side that is to play next is still highlighted. Also, unless searchTime or searchDepth is set, the chess program still keeps track of the clock time and uses it to determine how fast to make its moves.

-st or -searchTime minutes[:seconds] Tells the chess program to spend at most the given amount of time searching for each of its moves. Without this option, the chess program chooses its search time based on the number of moves and amount of time remaining until the next time control. Setting this option also sets clockMode to false.

-sd or -searchDepth number Tells the chess program to look ahead at most the given number of moves when searching for a move to make. Without this option, the chess program chooses its search depth based on the number of moves and amount of time remaining until the next time control. Setting this option also sets clockMode to false.

-thinking or -showThinking If this option is set, the chess program’s notion of the score and best line of play from the current position is displayed as it is thinking. The score indicates how many pawns ahead (or if negative, behind) the chess program thinks it is. In matches between two machines, the score is prefixed by ‘W’ or ‘B’ to indicate whether it is showing White’s thinking or Black’s.

-mm or -matchMode Automatically runs a game between two chess programs. If the loadGameFile or loadPositionFile option is set, XBoard starts the game with the given opening moves or the given position; otherwise, the game starts with the standard initial chess position. If the saveGameFile option is set, a move record for the match is appended to the specified file. If the savePositionFile option is set, the final position reached in the match is appended to the specified file. When the match is over, XBoard exits. Default: false.

-fcp or -firstChessProgram program Name of first chess program. In matches between two machines, this program plays Black. Default: ‘gnuchessx’.

-scp or -secondChessProgram program Name of second chess program, if needed. In matches between two machines, this program plays White; otherwise it is not started. Default: ‘gnuchessx’.

-fh or -firstHost host -sh or -secondHost host

Hosts on which the chess programs are to run. The default for each is ‘localhost’. If you specify another host, XBoard uses ‘rsh’ (1) to run the chess program there. (You can substitute a different remote shell program for rsh using the remoteShell option described below.)

-initString string

The string that is sent to initialize the chess program. Default:

     new
     beep
     random
     easy

Setting this option from the command line is tricky, because you must type in real newline characters, including one at the very end. In most shells you can do this by entering a ‘\’ character followed by a newline. It is easier to set the option from your ‘.Xdefaults’ file; in that case you can include the character sequence ‘\n’ in the string, and it will be converted to a newline.

If you change this option, don’t remove the ‘new’ and ‘beep’ commands. You can remove the ‘random’ command if you like; including it causes GNU Chess to randomize its move selection slightly so that it doesn’t play the same moves in every game. (Even without ‘random’, GNU Chess randomizes its choice of moves from its opening book.) You can also remove ‘easy’ if you like; including it toggles easy mode off, causing GNU Chess to think on your time. That is, if ‘easy’ is included in the initString, GNU Chess thinks on your time; if not, it does not. (Yes, this does seem backwards, doesn’t it!) You can also try adding other commands to the initString; see the GNU Chess documentation for details.

-whiteString string -blackString string

These options control what is sent when the Machine White and Machine Black buttons are selected. This is mostly for compatibility with obsolete versions of GNU Chess.

3.2 Connecting to and using ICS

-ics or -internetChessServerMode Connect with an Internet Chess Server to play chess against its other users, observe games they are playing, or review games that have recently finished. Default: false. To use XBoard in ICS mode, run it in the foreground, and use the terminal you started it from to type commands and receive text responses from the chess server. See Chapter 5 [ICS], page 25. See Section 4.1 [File Menu], page 16. See Section 4.4 [Step Menu], page 20. See Section 4.2 [Mode Menu], page 17.

-icslogon or -internetChessServerLogonScript file-name Whenever XBoard connects to the Internet Chess Server, if it finds a file with the name given in this option, it feeds the file’s contents to the ICS as commands. The default file name is ‘.icsrc’. Usually the first two lines of the file should be your ICS user name and password. The file can be either in $CHESSDIR, in XBoard’s working directory if CHESSDIR is not set, or in your home directory.

-msLoginDelay delay If you experience trouble logging on to an ICS when using the -icslogon option, inserting some delay between characters of the logon script may help. This option adds delay milliseconds of delay between characters. Good values to try are 100 and 250.

-icsinput or -internetChessServerInputBox If this option is True in ICS mode, XBoard creates an extra window that you can use for typing in ICS commands. The input box is especially useful if you want to type in something long or do some editing on your input, because output from ICS doesn’t get mixed in with your typing as it would in the main terminal window.

-autocomm or -autoComment If autoComment is True, any remarks made on ICS while you are observing or playing a game are recorded as a comment on the current move. This includes remarks made with the ICS commands say, tell, whisper, and kibitz. Limitation: remarks that you type yourself are not recognized; XBoard scans only the output from ICS, not the input you type to it. Default: False.

-autoflag or -autoCallFlag If autoCallFlag is true and your opponent runs out of time before you do, XBoard will automatically call his flag, claiming a win on time (or a draw if you do not have mating material). Default: false.

-autobs or -autoObserve If autoObserve is true and you add a player to your gnotify list on ICS, XBoard will automatically observe all of that player’s games, unless you are doing something else (such as observing or playing a game of your own) when one starts. The games are displayed from the point of view of the player on your gnotify list; that is, his pawns move from the bottom of the window towards the top. Exceptions: If both players in a game are on your gnotify list, if your ICS highlight variable is set to 0, or if the ICS you are using does not properly support observing from Black’s point of view, you will see the game from White’s point of view. Default: false.

-moves or -getMoveList If getMoveList is True, whenever XBoard receives the first board of a new game (or a different game from the one it is currently displaying), it retrieves the list of past moves from the ICS. You can then review the moves with the ‘Forward’ and ‘Backward’ commands or save them with ‘Save Game’. Default: true. You might want to turn off this option if you are observing several blitz games at once, to keep from wasting time and network bandwidth fetching the move lists over and over.

-quiet or -quietPlay

If this option is true, XBoard will automatically issue an ICS

     set shout 0

command whenever you start a game and a

     set shout 1

command whenever you finish one. Thus, you will not be distracted by shouts from other ICS users while playing. Default: false.

-icshost or -internetChessServerHost host The Internet host name or address of the chess server to connect to when in ICS mode. Default: chessclub.com. See the file ‘ics-addresses’ in the XBoard source distribution for a list of other addresses to try. See also the output of the command

     finger chess@ics.onenet.net

If your site doesn’t have a working Internet name server, try specifying the host address in numeric form. The address of chessclub.com is ‘192.231.221.16’; that of ics.onenet.net is 164.58.253.10.

-icsport or -internetChessServerPort port-number The port number to use when connecting to a chess server in ICS mode. Default: 5000.

-gateway host-name If this option is set to a host name, XBoard uses ‘rsh’ (1) to run the ‘telnet’ (1) program on the given host to communicate with the Internet Chess Server instead of using its own internal implementation of the telnet protocol. You can substitute a different remote shell program for ‘rsh’ using the remoteShell option described below. See Chapter 6 [Firewalls], page 27.

-telnet or -useTelnet If this option is set to true, XBoard runs the ‘telnet’ (1) program to communicate with the Internet Chess Server. If the option is false (the default), XBoard opens a TCP socket and uses its own internal implementation of the telnet protocol to communicate with the ICS. See Chapter 6 [Firewalls], page 27.

-telnetProgram prog-name This option gives the name of the telnet program to be used with the gateway and useTelnet options. The default is ‘telnet’. The telnet program is invoked with the value of internetChessServerHost as its first argument and the value of internetChessServerPort as its second argument. See Chapter 6 [Firewalls], page 27.

-internetChessServerCommPort or -icscomm dev-name If this option is set, XBoard communicates with the ICS through the given character I/O device instead of opening a TCP connection. Use this option if your system does not have any kind of Internet connection itself (not even a SLIP or PPP connection), but you do have dialup access (or a hardwired terminal line) to an Internet service provider from which you can telnet to the ICS. The support for this option in XBoard is minimal. You need to set all communication parameters and tty modes before you enter XBoard. Use a script something like this:

     stty raw -echo 9600 > /dev/tty00
     xboard -ics -icscomm /dev/tty00

Here replace ‘/dev/tty00’ with the name of the device that your modem is connected to. You might have to add several more options to these stty commands. See the man pages for ‘stty’ (1) and tty (4) if you run into problems. Also, on many systems stty works on its standard input instead of standard output, so you have to use ‘<’ instead of ‘>’. If you are using linux, try starting with the script below. Change it as necessary for your installation.

     ####################################################
     #!/bin/sh -f
     # configure modem and fire up XBoard
     # configure modem
     (
       stty 2400 ; stty raw ; stty hupcl ; stty -clocal
       stty ignbrk ; stty ignpar ; stty ixon ; stty ixoff
       stty -iexten ; stty -echo
     ) < /dev/modem
     xboard -ics -icscomm /dev/modem
     ####################################################

After you start XBoard in this way, type whatever commands are necessary to dial out to your Internet provider and log in. Then telnet to ICS, using a command like

     telnet chessclub.com 5000

Important: See the paragraph below about extra echoes, in Chapter 9 [Limitations], page 34.

3.3 Load and Save options

-lgf or -loadGameFile file -lgi or -loadGameIndex index

If the loadGameFile option is set, XBoard loads the specified game file at startup. The file name ‘-’ specifies the standard input. If there is more than one game in the file, XBoard pops up a menu of the available games, with entries based on their PGN tags. If the loadGameIndex option is set to ‘N’, the menu is suppressed and the N th game found in the file is loaded immediately. The menu is also suppressed if matchMode is enabled or if the game file is a pipe; in these cases the first game in the file is loaded immediately. Use the ‘pxboard’ shell script provided with XBoard if you want to pipe in files containing multiple games and still see the menu.

-td or -timeDelay seconds Time delay between moves during ‘Load Game’. Fractional seconds are allowed; try ‘-td 0.4’. A time delay value of -1 tells XBoard not to step through game files automatically. Default: 1 second.

-sgf or -saveGameFile file If this option is set, XBoard appends a record of every game played to the specified file. The file name ‘-’ specifies the standard output.

-autosave or -autoSaveGames If this option is true, at the end of every game XBoard prompts you for a file name and appends a record of the game to the file you specify. Ignored if saveGameFile is set. Default: false.

-lpf or -loadPositionFile file -lpi or -loadPositionIndex index

If the loadPositionFile option is set, XBoard loads the specified position file at startup. The file name ‘-’ specifies the standard input. If the loadPositionIndex option is set to N, the Nth position found in the file is loaded; otherwise the first position is loaded.

-spf or -savePositionFile file If this option is set, XBoard appends the final position reached in every game played to the specified file. The file name ‘-’ specifies the standard output.

-oldsave or -oldSaveStyle If this option is false (the default), XBoard saves games in PGN (portable game notation) and positions in FEN (Forsythe-Edwards notation). If the option is true, a save style that is compatible with older versions of XBoard is used instead.

3.4 Look and Feel options

-display -geometry -iconic These are just the standard Xt options accepted by XBoard.

-bell or -ringBellAfterMoves If this option is true, XBoard alerts you by ringing the terminal bell after each of your opponent’s moves (or after every move if you are observing a game on the Internet Chess Server). The bell is not rung after moves you make or moves read from a saved game file. Default: false. If you turn on this option when using XBoard with the Internet Chess Server, you will probably want to give the

     set bell 0

command to the ICS, since otherwise the ICS will ring the bell itself after every move (not just yours). (The ‘.icsrc’ file is a good place for this, see Section 3.2 [ICS options], page 7)

-queen or -alwaysPromoteToQueen If this option is false (the default), XBoard brings up a dialog box whenever you move a pawn to the last rank, asking what piece you want to promote it to. If the option is true, your pawns are always promoted to queens. (Your opponent can still underpromote, however.)

-legal or -checkLegality If this option is true (the default), XBoard checks whether moves you try to make with the mouse are legal, and refuses to let you make an illegal move. If the option is false, you can make any move. If GNU Chess, Crafty, or the ICS is active, they will still reject illegal moves. Turning off this option is useful if you are playing a chess variant with rules that XBoard does not understand.

-size or -boardSize (Large | Medium | Small | Tiny | n1,n2,n3,n4,n5,n6)

Determines how large the board will be. The Large board uses 80x80 pieces, Medium 64x64, Small 40x40, and Tiny 21x21. Piece bitmaps of all these sizes are built into XBoard. The default depends on the size of your screen; it is approximately the largest size that will fit without clipping.

You can select other sizes or vary other layout parameters by providing a list of commaseparated values (with no spaces) as the argument. The n1 value gives the piece size, n2 the width of the black border between squares, n3 the preferred pixel size for the clockFont, n4 the preferred pixel size for the coordFont, n5 the smallLayout flag (0 or 1), and n6 the tinyLayout flag (0 or 1). If smallLayout is 1 and titleInWindow is True, the window layout is rearranged to make more room for the title. If tinyLayout is 1, the labels on the menu bar are abbreviated to one character each, the buttons in the button bar are made narrower, and a smaller default font is used. You do not need to provide all the values; for any you omit from the end of the list, defaults are taken from the nearest built-in size.

-coords or -showCoords If this option is true, XBoard displays algebraic coordinates along the board’s left and bottom edges. The default is false. The coordFont option specifies what font to use.

-flip or -flipView If you are playing a game on the ICS, the board is always oriented at the start of the game so that your pawns move from the bottom of the window towards the top. Otherwise, the starting orientation is determined by the flipView option; if it is false (the default), White’s pawns move from bottom to top at the start of each game; if it is true, Black’s pawns move from bottom to top.

-title or -titleInWindow If this option is true, XBoard displays player names (for ICS games) and game file names (for ‘Load Game’) inside its main window. If the option is false (the default), this information is displayed only in the window banner. You probably won’t want to set this option unless the information is not showing up in the banner, as happens with a few X window managers.

-mono or -monoMode Determines whether XBoard displays its pieces and squares with two colors (true) or four (false). You shouldn’t have to specify monoMode; XBoard will determine if it is necessary.

-borderXoffset -borderYoffset

These options are part of a kludge that helps position the Comment and Edit Comment windows in the same place on the screen when they are closed and reopened. They should be set equal to the width and height of the borders that your X window manager adds to windows when it displays them. The defaults are correct for ‘tvtwm’ (1).

-flashCount count -flashRate rate

These options control whether pieces flash after landing on their destination square. For example, to have pieces flash 3 times at a rate of 5 flashes/second, you would say:

           xboard -flashCount 3 -flashRate 5  ... (other options) ...

Or, just put them in your .Xdefaults file so you don’t have to give the command line options every time:

           xboard*flashCount: 3
           xboard*flashRate: 5

-clickClick When "Click-Click" mode is enabled (-clickClick True), you move pieces by first clicking on the piece, then clicking on the destination square. When "Click-Click" mode is disabled (the default) you move pieces by clicking on the piece, "dragging" it to the destination square, and releasing the mouse button to drop it.

-clockFont The font used for the clocks. If the option value is a pattern that does not specify the font size, XBoard tries to choose an appropriate font for the board size being used. Default: -*-helvetica-bold-r-normal–*-*-*-*-*-*-*-*.

-coordFont The font used for rank and file coordinate labels if showCoords is true. If the option value is a pattern that does not specify the font size, XBoard tries to choose an appropriate font for the board size being used. Default: -*-helvetica-bold-r-normal–*- *-*-*-*-*-*-*.

-font The font used for popup dialogs, menus, comments, etc. If tinyLayout is true (e.g., if "-size Tiny" is specified), the default is -*-helvetica-medium-r-normal–11-*-*-*-*-*-*-*. Otherwise, the default is -*-helvetica-medium-r-normal–14-*-*-*-*-*-*-*.

-bm or -bitmapDirectory By default, XBoard uses a set of compiled-in bitmaps for its pieces, icons, and menu checkmark. If the bitmapDirectory option is set at runtime, bitmaps are taken from files in the specified directory. If any bitmap file is missing or unusable, XBoard looks for a built-in bitmap of the required type and size instead. Files in the bitmapDirectory must be named as follows: The first character of a piece bitmap name gives the piece it represents (‘p’, ‘n’, ‘b’, ‘r’, ‘q’, or ‘k’), the next characters give the size in pixels, the following character indicates whether the piece is solid or outline (‘s’ or ‘o’), and the extension is ‘.bm’. For example, a solid 80x80 knight would be named ‘n80s.bm’. The outline bitmaps are used only in monochrome mode. The icons are named ‘icon_white.bm’ and ‘icon_black.bm’, and the menu checkmark is

named ‘checkmark.bm’.

Two sets of bitmaps are distributed with XBoard. Those in the directory ‘bitmaps’ are normally compiled in as the default. Those in the ‘bitmaps.xchess’ directory can be selected at runtime with the bitmapDirectory option. If you want to compile in the latter set as the default, rename the ‘bitmaps’ directory to ‘bitmaps.fselch’ and the ‘bitmaps.xchess’ directory to ‘bitmaps’; then recompile XBoard. If you want to add another compiled-in size, edit the ‘bitmaps.h’ file in the bitmaps directory, and optionally edit ‘xboard.h’ to associate a name and default layout parameters with your new size.

-xpm directory Instead of bitmaps, XBoard can use X pixmaps as its piece images. If the xpmDirectory option is set, pixmaps are taken from files in the specified directory. The directory may contain multiple sizes of pieces. XBoard will choose the closest based either on screen size or on the size option. You can also specify this option in your .Xdefaults file like this:

          xboard*xpmDirectory: ~/pieces

For more information about XPM pieces, see Section 1.3 [zic2xpm], page 2 -colorizeMessages or -colorize

Setting colorizeMessages to True tells XBoard to colorize the messages received from the ICS. Colorization works only if your xterm supports ISO 6429 escape sequences for changing text colors.

-colorShout foreground,background,bold -colorSShout foreground,background,bold -colorChannel1 foreground,background,bold -colorChannel foreground,background,bold -colorKibitz foreground,background,bold -colorTell foreground,background,bold -colorChallege foreground,background,bold -colorNormal foreground,background,bold

These options set the colors used when colorizing ICS messages. All ICS messages are grouped into one of these categories: shout, sshout, channel 1, other channel, kibitz, tell, challenge, normal. Each foreground or background argument can be one of the following: black, red, green, yellow, blue, magenta, cyan, white, or default. Here “default” means the default foreground or background color of your xterm. Bold can be 1 or 0. If background is omitted, “default” is assumed; if bold is omitted, 0 is assumed. Here is an example of how to set the colors in your ‘.Xdefaults’ file. The colors shown here are the default values; you will get them if you turn -colorize on without specifying your own colors.

         xboard*colorizeMessages: true
         xboard*colorShout: green
         xboard*colorSShout: green, black, 1
         xboard*colorChannel1: cyan
         xboard*colorChannel: cyan, black, 1
         xboard*colorKibitz: magenta, black, 1
         xboard*colorTell: yellow, black, 1
         xboard*colorChallenge: red, black, 1
         xboard*colorNormal: default

-whitePieceColor -blackPieceColor -lightSquareColor -darkSquareColor

Colors to use for the pieces and squares. Defaults:

     -whitePieceColor    #FFFFCC
     -blackPieceColor    #202020
     -lightSquareColor   #C8C365
     -darkSquareColor    #77A26D

On a grayscale monitor you might prefer:

     -whitePieceColor    gray100
     -blackPieceColor    gray0

     -lightSquareColor   gray80
     -darkSquareColor    gray60

3.5 Miscellaneous

-ncp or -noChessProgram If this option is true, XBoard acts as a passive chessboard; it does not start a chess program at all. Turning on this option also turns off clockMode. Default: false.

-debug or -debugMode Turns on debugging printout.

-rsh or -remoteShell shell-name Name of the command used to run programs remotely. The default is ‘rsh’ or ‘remsh’, determined when XBoard is configured and compiled.

-ruser or -remoteUser user-name User name on the remote system when running programs with the remoteShell. The default is your local user name.

4 Menus, buttons and keys

To move a piece, drag it with the left mouse button. To drop a new piece on a square (when applicable), press the middle or right mouse button over the square and select from the popup menu.

All other XBoard commands are available from the menu bar. The most frequently used commands also have shortcut keys or on-screen buttons.

4.1 Accessing external games and positions

Reset Resets XBoard and the chess program to the beginning of a new chess game. The ‘r’ key is a keyboard equivalent. In Internet Chess Server mode, clears the current state of XBoard, then resynchronizes with ICS by sending a refresh command. If you want to stop playing, observing, or examining a game on ICS, use an appropriate command from the Action menu, not ‘Reset’. See Section 4.3 [Action Menu], page 19.

Load Game Plays a game from a record file. The ‘g’ key is a keyboard equivalent. A popup dialog prompts you for the file name. If the file contains more than one game, a second popup dialog displays a list of games (with information drawn from their PGN tags, if any), and you can select the one you want. Alternatively, you can load the Nth game in the file directly, by typing the number ‘N’ after the file name, separated by a space. The game file parser will accept PGN (portable game notation), or in fact almost any file that contains moves in algebraic notation. Notation of the form ‘P@f7’ is accepted for piece-drops in bughouse games; this is a nonstandard extension to PGN. If the file includes a PGN position (FEN tag), or an XBoard position diagram bracketed by ‘[--’ and ‘--]’ before the first move, the game starts from that position. Text enclosed in parentheses, square brackets, or curly braces is assumed to be commentary and is displayed in a pop-up window. Any other text in the file is ignored. PGN variations (enclosed in parentheses) are treated as comments; XBoard is not able to walk variation trees.

Load Next Game Loads the next game from the last game record file you loaded. The shifted ‘N’ key is a keyboard equivalent. Not available if the last game was loaded from a pipe.

Load Previous Game Loads the previous game from the last game record file you loaded. The shifted ‘P’ key is a keyboard equivalent. Not available if the last game was loaded from a pipe.

Reload Same Game Reloads the last game you loaded.

Load Position Sets up a position from a position file. A popup dialog prompts you for the file name. If the file contains more than one saved position, and you want to load the Nth one, type the number N after the file name, separated by a space. Position files must be in FEN (Forsythe-Edwards notation), or in the format that the Save Position command writes when oldSaveStyle is turned on.

Save Game Appends a record of the current game to a file. A popup dialog prompts you for the file name. If the game did not begin with the standard starting position, the game file includes the starting position used. Games are saved in the PGN (portable game notation) format, unless the oldSaveStyle option is true, in which case they are saved in an older format that is specific to XBoard. Both formats are human-readable, and both can be read back by the ‘Load Game’ command. Notation of the form ‘P@f7’ is accepted for piece-drops in bughouse games; this is a nonstandard extension to PGN.

Save Position Appends a diagram of the current position to a file. A popup dialog prompts you for the file name. Positions are saved in FEN (Forsythe-Edwards notation) format unless the oldSaveStyle option is true, in which case they are saved in an older, humanreadable format that is specific to XBoard. Both formats can be read back by the

‘Load Position’ command.

Mail Move Reload CMail Message

See Chapter 7 [CMail], page 29.

Exit Exits from XBoard. The shifted ‘Q’ key is a keyboard equivalent.

4.2 Selecting XBoard’s mode

Machine White Forces the chess program to play White.

Machine Black Forces the chess program to play Black.

Two Machines Plays a game between two computer programs.

Analysis Mode XBoard with Crafty only. For information on getting Crafty, see Section 1.2 [Crafty], page 2. XBoard tells Crafty to start analyzing the current game/position and shows you the analysis as you move pieces around. To setup a position to analyze, you do the following:

1. Select Edit Position from the Mode Menu

2. Setup the position (use the right and middle buttons to bring up the Black/White piece menus).

3. When you are finished, click on either the Black or White clock to tell XBoard which side moves first.

4. Select Analysis Mode from the Mode Menu to start the analysis.

Analyze File XBoard with Crafty only. For information on getting Crafty, see Section 1.2 [Crafty], page 2.

This option lets you load a game from a file (PGN, XBoard format, etc.) and analyze it. When you select this menu item, a popup window appears and asks for a filename to load. If the file contains multiple games, another popup appears that lets you select which game you wish to analyze. After a game is loaded, use the XBoard arrow buttons to step forwards/backwards through the game and watch Crafty’s analysis.

ICS Client ICS mode only. Takes XBoard out of the Edit Game or Edit Position state. While you are examining a game on the ICS, you can issue the ICS position-editing commands with the mouse. Do this with ‘ICS Client’ selected on the Mode menu, not ‘Edit Position’; the latter edits only your local copy of the position. To drop a new piece on a square, press mouse button 2 or 3 over the square. This brings up a menu of white pieces (button 2) or black pieces (button 3). Additional menu choices let you empty the square or clear the board. You cannot set the side to play or drag pieces to arbitrary squares while examining on ICS, however; the ICS permits only legal moves in this mode. If you are playing a bughouse game on the ICS, you can drop an offboard piece by pressing mouse button 2 or 3 over an empty square to bring up a piece menu. It makes no difference which button you use. A list of the offboard pieces each player has available is shown in the window title after the player’s name.

Edit Game Allows you to make moves for both Black and White, and to change moves after backing up with the ‘Backward’ command. The clocks do not run. In GNU Chess mode, the chess program continues to check moves for legality but does not participate in the game. You can bring the chess program into the game by

selecting ‘Machine White’, ‘Machine Black’, or ‘Two Machines’.

In ICS mode, the moves are not sent to the ICS: ‘Edit Game’ takes XBoard out of ICS Client mode and lets you edit games locally. If you want to edit games on ICS in a way that other ICS users can see, use the ICS examine command or start an ICS match against yourself.

Edit Position Lets you set up an arbitrary board position. Use mouse button 1 to drag pieces to new squares, or to delete a piece by dragging it off the board or dragging an empty square on top of it. To drop a new piece on a square, press mouse button 2 or 3 over the square. This brings up a menu of white pieces (button 2) or black pieces (button 3). Additional menu choices let you empty the square or clear the board. You can set the side to play next by clicking on the White or Black indicator at the top of the screen. Selecting ‘Edit Position’ causes XBoard to discard all remembered moves in the current game. In ICS mode, changes made to the position by ‘Edit Position’ are not sent to the ICS: ‘Edit Position’ takes XBoard out of ‘ICS Client’ mode and lets you edit positions locally. If you want to edit positions on ICS in a way that other ICS users can see, use the ‘ICS examine’ command, or start an ICS match against yourself. (See also the ICS Client topic above.)

Show Game List Shows or hides the list of games generated by the last ‘Load Game’ command.

Edit Tags Lets you edit the PGN (portable game notation) tags for the current game. After

editing, the tags must still conform to the PGN tag syntax:

     <tag-section> ::= <tag-pair> <tag-section>
                             <empty>
     <tag-pair> ::= [ <tag-name> <tag-value> ]
     <tag-name> ::= <identifier>
     <tag-value> ::= <string>

See the PGN Standard for full details. Here is an example:

     [Event "Portoroz Interzonal"]
     [Site "Portoroz, Yugoslavia"]
     [Date "1958.08.16"]
     [Round "8"]
     [White "Robert J. Fischer"]
     [Black "Bent Larsen"]
     [Result "1-0"]

Any characters that do not match this syntax are silently ignored. Note that the PGN standard requires all games to have at least the seven tags shown above. Any that you omit will be filled in by XBoard with ‘?’ (unknown value), or ‘-’ (inapplicable value).

Edit Comment Adds or modifies a comment on the current position. Comments are saved by ‘Save Game’ and are displayed by ‘Load Game’, ‘Forward’, and ‘Backward’.

ICS Input Box Shows or hides the ICS input box. See the internetChessServerInputBox option. ICS mode only.

Pause Pauses updates to the board, and if you are playing against a chess program, also pauses your clock. To continue, select ‘Pause’ again, and the display will automatically update to the latest position. The ‘P’ button and keyboard ‘p’ key are equivalents. If you select Pause when you are playing against a chess program and it is not your move, the chess program’s clock will continue to run and it will eventually make a move, at which point both clocks will stop. Since board updates are paused, however, you will not see the move until you exit from Pause mode (or select Forward). This behavior is meant to simulate adjournment with a sealed move. If you select Pause while you are in examine mode on ICS, you can step backward and forward in the current history of the examined game without affecting the other observers and examiners. Select Pause again to reconnect yourself to the current state of the game on ICS. If you select ‘Pause’ while you are loading a game, the game stops loading. You can load more moves manually by selecting ‘Forward’, or resume automatic loading by selecting ‘Pause’ again.

4.3 Talking to GNU chess or ICS opponents

Accept Accepts a pending match offer. If there is more than one offer pending, you will have to type in a more specific command instead of using this menu choice. ICS mode only.

Decline Declines a pending offer (match, draw, adjourn, etc.). If there is more than one offer pending, you will have to type in a more specific command instead of using this menu choice. ICS mode only.

Call Flag Calls your opponent’s flag, claiming a win on time, or claiming a draw if you are both out of time. You can also call your opponent’s flag by clicking on his clock or by pressing the keyboard ‘t’ key.

Draw Offers a draw to your opponent, accepts a pending draw offer from your opponent, or claims a draw by repetition or the 50-move rule, as appropriate. The ‘d’ key is a keyboard equivalent. Not available in GNU Chess mode.

Adjourn Asks your opponent to agree to adjourning the current game, or agrees to a pending adjournment offer from your opponent. ICS mode only.

Abort Asks your opponent to agree to aborting the current game, or agrees to a pending abort offer from your opponent. An aborted game ends immediately without affecting either player’s rating.

Resign Resigns the game to your opponent. The shifted ‘R’ key is a keyboard equivalent.

Stop Observing Ends your participation in observing a game, by issuing the ICS observe command with no arguments. ICS mode only.

Stop Examining Ends your participation in examining a game, by issuing the ICS unexamine command. ICS mode only.

4.4 Controlling the game

Backward Steps backward through a series of remembered moves. The ‘[<]’ button and the ‘b’ key are equivalents. In addition, pressing the Control key steps back one move, and releasing it steps forward again. In most modes, ‘Backward’ only lets you look back at old positions; it does not retract moves. This is the case if you are playing against a chess program, playing or observing a game on the ICS, or loading a game. If you select ‘Backward’ in any of these situations, you will not be allowed to make a different move. Use ‘Retract Move’ or ‘Edit Game’ if you want to change past moves. If you are examining a game on ICS, the behavior of ‘Backward’ depends on whether XBoard is in Pause mode. If Pause mode is off, ‘Backward’ issues the ICS backward command, which backs up everyone’s view of the game and allows you to make a different move. If Pause mode is on, ‘Backward’ only backs up your local view.

Forward Steps forward through a series of remembered moves (undoing the effect of ‘Backward’) or forward through a game file. The ‘[>]’ button and the ‘f’ key are equivalents. If you are examining a game on ICS, the behavior of Forward depends on whether XBoard is in Pause mode. If Pause mode is off, ‘Forward’ issues the ICS forward command, which moves everyone’s view of the game forward along the current line. If Pause mode is on, ‘Forward’ only moves your local view forward, and it will not go past the position that the game was in when you paused.

Back to Start Jumps backward to the first remembered position in the game. The ‘[<<]’ button and the shifted ‘B’ key are equivalents. In most modes, Back to Start only lets you look back at old positions; it does not retract moves. This is the case if you are playing against GNU chess, playing or observing a game on the ICS, or loading a game. If you select ‘Back to Start’ in any of these situations, you will not be allowed to make different moves. Use ‘Retract Move’ or ‘Edit Game’ if you want to change past moves; or use Reset to start a new game. If you are examining a game on ICS, the behavior of ‘Back to Start’ depends on whether XBoard is in Pause mode. If Pause mode is off, ‘Back to Start’ issues the ICS ‘backward 999999’ command, which backs up everyone’s view of the game to the start and allows you to make different moves. If Pause mode is on, ‘Back to Start’ only backs up your local view.

Forward to End Jumps forward to the last remembered position in the game. The ‘[>>]’ button and the shifted ‘F’ key are equivalents. If you are examining a game on ICS, the behavior of ‘Forward to End’ depends on whether XBoard is in Pause mode. If Pause mode is off, ‘Forward to End’ issues the ICS ‘forward 999999’ command, which moves everyone’s view of the game forward to the end of the current line. If Pause mode is on, ‘Forward to End’ only moves your local view forward, and it will not go past the position that the game was in when you paused.

Revert If you are examining a game on ICS and Pause mode is off, issues the ICS command

‘revert’.

Truncate Game Discards all remembered moves of the game beyond the current position. Puts XBoard into ‘Edit Game’ mode if it was not there already.

Move Now Forces the chess program to move immediately. GNU Chess mode only.

Retract Move Retracts your last move. In GNU Chess mode, you can do this only after the chess program has replied to your move; if the chess program is still thinking, use ‘Move Now’ first. In ICS mode, ‘Retract Move’ issues the command ‘takeback 1’ or ‘takeback 2’ depending on whether it is your opponent’s move or yours.

4.5 User Preferences

Always Queen Toggles the alwaysPromoteToQueen option. See Section 3.4 [User interface options], page 11.

Auto Comment Toggles the autoComment option. See Section 3.2 [ICS options], page 7.

Auto Flag Toggles the autoCallFlag option: XBoard will automatically claim a win for you, if your opponents flag falls on ICS. See Section 3.2 [ICS options], page 7.

Auto Observe Toggles the autoObserve option. See Section 3.2 [ICS options], page 7.

Auto Save Toggles the autoSaveGames option. Disabled if the saveGameFile option is set, as in that case all games are saved to the specified file. See Section 3.3 [I/O options], page 10.

Bell Toggles the ringBellAfterMoves option. See Section 3.4 [User interface options], page 11.

Flip View Inverts your view of the chess board for the duration of the current game. Starting a new game returns the board to normal. If you are playing a game on the ICS, the board is always oriented at the start of the game so that your pawns move from the bottom of the window towards the top. Otherwise, the starting orientation is determined by the flipView command line option; if it is false (the default), White’s pawns move from bottom to top at the start of each game; if it is true, Black’s pawns move from bottom to top. See Section 3.4 [User interface options], page 11.

Get Move List Toggles the getMoveList option. See Section 3.2 [ICS options], page 7. When you turn this option on from the menu, XBoard immediately fetches the move list of the current game (if any).

Old Save Style Toggles the oldSaveStyle option. See Section 3.3 [I/O options], page 10.

Periodic Updates If you are using the "XBoard-supported" version of Crafty (see See Section 1.2 [Crafty], page 2) then you can decide whether you want the Analysis Window to be updated periodically or not. If this option is unchecked (or if you are not using the "XBoardsupported" Crafty), the the analysis window will only be updated when the analysis changes. If this option is checked, the Analysis Window will be updated every two seconds.

Quiet Play Toggles the quietPlay option. See Section 3.2 [ICS options], page 7.

Show Coords Toggles the showCoords option. See Section 3.4 [User interface options], page 11.

Show Thinking Toggles the showThinking option. See Section 3.1 [GNU Chess options], page 5.

4.6 Getting help

Info XBoard Displays this info file in a new window. For this feature to work, you must have the GNU info program installed on your system, and the file ‘xboard.info’ must either be present in the current working directory, or have been installed by the ‘make install’ command when you built XBoard.

Man XBoard Displays the XBoard man page in a new window. For this feature to work, the file ‘xboard.6’ must have been installed by the ‘make install’ command when you built XBoard, and the directory it was placed in must be on the search path for your system’s ‘man’ command.

Hint Displays a move hint from GNU Chess. GNU Chess mode only.

Book Displays a list of possible moves from GNU Chess’s opening book. The first column gives moves, the second column gives one possible response for each move, and the third column shows the number of lines in the book that include the move from the first column. If you select this option and nothing happens, GNU Chess is out of its book. GNU Chess mode only.

About XBoard Shows the current XBoard version number.

4.7 Other shortcut keys

Iconize Pressing the ‘i’ or ‘c’ key iconizes XBoard. The graphical icon displays a white knight if it is white’s move, or a black knight, if it is Black’s move. If your X window manager displays only text icons, not graphical ones, check its documentation; there is probably a way to enable graphical icons. If you are running the Motif window manager ‘mwm’ (1), add these lines to your ‘.Xdefaults’ file and restart mwm:

     Mwm*iconDecoration: activelabel label image
     Mwm*XBoard*iconImageBackground: White
     Mwm*XBoard*iconImageForeground: Black

The first line above enables graphical icons in mwm; you don’t need it if you already have them. The next two lines force the white knights to come out white and the black knights black. Unfortunately these resources can’t be set from inside XBoard; you have to set them in your ‘.Xdefaults’ file.

You can add or remove shortcut keys using the X resources form.translations. Here is an example of what would go in your ‘.Xdefaults’ file:

XBoard*form.translations: Shift<Key>?: AboutGameProc() \n\
<Key>y: AcceptProc() \n\
<Key>n: DeclineProc() \n\
<Key>i: NothingProc()

Binding a key to NothingProc makes it do nothing, thus removing it as a shortcut key. The XBoard commands that can be bound to keys are:

AbortProc, AboutGameProc, AboutProc, AcceptProc, AdjournProc,
AlwaysQueenProc, AnalyzeGameProc, AnalyzeModeProc, AutobsProc,
AutoflagProc, AutosaveProc, BackwardProc, BellProc, BookProc,
CallFlagProc, DebugProc, DeclineProc, DrawProc, EditCommentProc,
EditGameProc, EditPositionProc, EditTagsProc, FlipViewProc, ForwardProc,
GetMovesProc, HintProc, Iconify, IcsClientProc, InfoProc, LoadGameProc,
LoadNextGameProc, LoadPositionProc, LoadPrevGameProc, MachineBlackProc,
MachineWhiteProc, MailMoveProc, ManProc, MoveNowProc, NothingProc,

OldSaveStyleProc, PauseProc, PeriodicUpdatesProc, QuietPlayProc,
QuitProc, ReloadCmailMsgProc, ReloadGameProc, ResetProc, ResignProc,
RetractMoveProc, RevertProc, SaveGameProc, SavePositionProc,
ShowCoordsProc, ShowGameListProc, ShowThinkingProc, StopExaminingProc,
StopObservingProc, ToEndProc, ToStartProc, TruncateGameProc,
TwoMachinesProc.

5 Using XBoard with an Internet Chess Server (ICS)

An Internet Chess Server, or ICS, is a place on the Internet where people can get together to play chess, watch other people’s games, or just chat. You can use either telnet or a client program like XBoard to connect to the server. ICS is getting more and more important for chess players: There are thousands of registered users on the different ICS hosts, and it is not unusual to meet 200 on both ICC and FICS. The number is increasing rapidly.

Most people can just type

     xboard -ics

to start XBoard as an ICS client. Invoking XBoard in this way connects you to the Internet Chess Club (ICC), a commercial ICS. You can log in there as a guest even if you do not have a paid account. To connect to the largest Free ICS (FICS), use the command

     xboard -ics -icshost ics.onenet.net

instead, or substitute a different host name to connect to your favorite ICS. The ‘ics-addresses’ in the XBoard distribution includes a list of ICS hosts, and the file ‘ics-info’ is a longer introduction to ICS. For a full description of command-line options that control the connection to ICS and change the default values of ICS options, see Section 3.2 [ICS options], page 7.

While you are running XBoard as an ICS client, you use the terminal window that you started XBoard from as a place to type in commands and read information that is not available on the chessboard.

The first time you need to use the terminal is to enter your login name and password, if you are a registered player. (You don’t need to do this manually; the icsLogon option can do it for you. see Section 3.2 [ICS options], page 7) If you are not registered, enter any name. If someone has already registered under that name, you’ll be asked for a password; just hit return and try again. Or on ICC, you can enter ‘g’ as your name, and ICC will pick a unique (but boring) name for you.

Some useful ICS commands include

help <topic>

to get help on the given <topic>. To get a list of possible topics type help without topic. Try the help command before you ask other people on the server for help. For example help register tells you how to become a registered ICS player.

who <flags>

to see a list of people who are logged on. Administrators (people you should talk to if you have a problem) are marked with the character ‘*’, an asterisk. The <flags> allow you to display only selected players: For example, who of shows a list of players who are interested in playing but do not have an opponent.

games to see what games are being played

match <player> [<mins>] [<inc>]

to challenge another player to a game. Both opponents get <mins> minutes for the game, and <inc> seconds will be added after each move. If another player challenges you, the server asks if you want to accept the challenge; use the accept or decline commands to answer.

accept

decline to accept or decline another player’s offer. The offer may be to start a new game, or to agree to a draw, adjourn or abort the current game. See Section 4.3 [Action Menu], page 19. If you have more than one pending offer (for example, if more than one player is challenging you, or if your opponent offers both a draw and to adjourn the game), you have to supply additional information, by typing something like accept <player>,

          accept draw, or draw.
draw
adjourn

abort asks your opponent to terminate a game by mutual agreement. Adjourned games can be continued later. Your opponent can either decline your offer or accept it (by typing the same command or typing accept). In some cases these commands work immediately, without asking your opponent to agree. For example, you can abort the game unilaterally if your opponent is out of time, and you can claim a draw by repetition or the 50-move rule if available simply by typing draw.

finger <player>

to get informations about the given <player>. (Default: yourself.)

vars to get a list of personal settings

set <var> <value>

to modify these settings

observe <player>

to observe an ongoing game of the given <player>.

examine

oldmoves to review a recently completed game

Some special XBoard features are activated when you are in examine mode on ICS. See the descriptions of the menu commands ‘Forward’, ‘Backward’, ‘Pause’, ‘ICS Client’, and ‘Stop Examining’ on the Section 4.4 [Step Menu], page 20, Section 4.2 [Mode Menu], page 17, and Section 4.5 [Options Menu], page 21.

6 Connecting to the ICS through a firewall

By default, XBoard communicates with an Internet Chess Server by opening a TCP socket directly from the machine it is running on to the ICS. If there is a firewall between your machine and the ICS, this won’t work. Here are some recipes for getting around common kinds of firewalls using special options to XBoard. Important: See the paragraph in the below about extra echoes, in Chapter 9 [Limitations], page 34.

Suppose that you can’t telnet directly to ICS, but you can telnet to a firewall host, log in, and then telnet from there to ICS. Let’s say the firewall is called ‘fire.wall.com’. Set command-line options as follows:

     xboard -ics -icshost fire.wall.com -icsport 23

Or in your ‘.Xdefaults’ file:

     XBoard*internetChessServerHost: fire.wall.com
     XBoard*internetChessServerPort: 23

Then when you run XBoard in ICS mode, you will be prompted to log in to the firewall host. (This works because port 23 is the standard telnet login service.) Do so, then telnet to ICS, using a command like ‘telnet chessclub.com 5000’, or whatever command the firewall provides for telnetting to port 5000.

If your firewall lets you telnet (or rlogin) to remote hosts, but doesn’t let you telnet to port 5000, you will have to find some other host outside the firewall that does let you do this, and hop through it. For instance, suppose you have an account at ‘foo.edu’. Follow the recipe above, but

instead of typing ‘telnet chessclub.com 5000’ to the firewall, type ‘telnet foo.edu’ (or ‘rlogin
foo.edu’), log in there, and then type ‘telnet chessclub.com 5000’.

Exception: chessclub.com itself lets you connect to the chess server on the default telnet port (23), which is what you get if you don’t specify a port to the telnet program. But the other chess servers don’t allow this.

Suppose that you can’t telnet directly to ICS, but you can use rsh to run programs on a firewall host, and that host can telnet to ICS. Let’s say the firewall is called ‘rsh.wall.com’. Set commandline options as follows:

     xboard -ics -gateway rsh.wall.com -icshost chessclub.com

Or in your ‘.Xdefaults’ file:

     XBoard*gateway: rsh.wall.com
     XBoard*internetChessServerHost: chessclub.com

Then when you run XBoard in ICS mode, it will connect to the ICS by using ‘rsh’ to run the

command ‘telnet chessclub.com 5000’ on host ‘rsh.wall.com’.

Suppose that you can telnet anywhere you want, but you have to run a special program called ‘ptelnet’ to do so.

First, we’ll consider the easy case, in which ‘ptelnet chessclub.com 5000’ gets you to the chess server. In this case set command line options as follows:

     xboard -ics -telnet -telnetProgram ptelnet

Or in your ‘.Xdefaults’ file:

     XBoard*useTelnet: true
     XBoard*telnetProgram: ptelnet

Then when you run XBoard in ICS mode, it will issue the command ‘ptelnet chessclub.com 5000’ to connect to the ICS.

Next, suppose that ‘ptelnet chessclub.com 5000’ doesn’t work; that is, your ‘ptelnet’ program doesn’t let you connect to alternative ports. In this case, you will have to find some other host outside the firewall that does let you do this, and hop through it. For instance, suppose you have an account at ‘foo.edu’. Set command line options as follows:

     xboard -ics -telnet -telnetProgram ptelnet -icshost foo.edu -icsport ""

Or in your ‘.Xdefaults’ file:

     XBoard*useTelnet: true
     XBoard*telnetProgram: ptelnet
     XBoard*internetChessServerHost: foo.edu
     XBoard*internetChessServerPort:

Then when you run XBoard in ICS mode, it will issue the command ‘ptelnet foo.edu’ to connect to your account at ‘foo.edu’. Log in there, then type ‘telnet chessclub.com 5000’.

7 Using XBoard for electronic correspondence chess

The ‘cmail’ program will help you play chess by email with opponents of your choice using XBoard as an interface.

You will usually run ‘cmail’ without giving any options.

7.1 Invoking CMail.

-h Displays ‘cmail’ usage information.

-c Shows the conditions of the GNU General Public License. See [Copying], page 37.

-w Shows the warranty notice of the GNU General Public License. See [Copying], page 37.

-v -xv Provides or inhibits verbose output from ‘cmail’ and XBoard, useful for debugging.

The -xv form also inhibits the cmail introduction message.

-mail -xmail Invokes or inhibits the sending of a mail message containing the move.

-xboard -xxboard Invokes or inhibits the running of XBoard on the game file.

-reuse

-xreuse Invokes or inhibits the reuse of an existing XBoard to display the current game.

-remail Resends the last mail message for that game. This inhibits running XBoard.

-game <name> The name of the game to be processed.

-wgames <number> -bgames <number> -games <number>

Number of games to start as White, as Black or in total. Default is 1 as white and none as black. If only one color is specified then none of the other color is assumed. If no color is specified then equal numbers of White and Black games are started, with the extra game being as White if an odd number of total games is specified.

-me <short name> -opp <short name>

A one-word alias for yourself or your opponent.

-wname <full name> -bname <full name> -name <full name> -oppname <full name>

The full name of White, Black, yourself or your opponent.

-wna <net address> -bna <net address> -na <net address> -oppna <net address>

The email address of White, Black, yourself or your opponent.

-dir <directory> The directory in which ‘cmail’ keeps its files. This defaults to the environment variable

$CMAIL_DIR or failing that, $CHESSDIR, ‘$HOME/Chess’ or ‘~/Chess’. It will be created

if it does not exist.

-arcdir <directory> The directory in which ‘cmail’ archives completed games. Defaults to the environment variable $CMAIL_ARCDIR or, in its absence, the same directory as cmail keeps its working files (above).

-mailprog <mail program> The program used by cmail to send email messages. This defaults to the environment

variable $CMAIL_MAILPROG or failing that ‘/usr/ucb/Mail’, ‘/usr/ucb/mail’ or ‘Mail’.

You will need to set this variable if none of the above paths fit your system.

-gamesFile <file> A file containing a list of games with email addresses. This defaults to the environment variable $CMAIL_GAMES or failing that ‘.cmailgames’.

-aliasesFile <file> A file containing one or more aliases for a set of email addresses. This defaults to the environment variable $CMAIL_ALIASES or failing that ‘.cmailaliases’.

-logFile <file> A file in which to dump verbose debugging messages that are invoked with the ‘-v’ option.

-event <event> The PGN Event tag (default ‘Email correspondence game’).

-site <site> The PGN Site tag (default ‘NET’).

-round <round> The PGN Round tag (default ‘-’, not applicable).

-mode <mode> The PGN Mode tag (default ‘EM’, Electronic Mail).

OTHER OPTIONS Any unrecognized flags will be passed to XBoard. Those most relevant for use with

Icmail\fP are ‘-timeDelay’, ‘-noChessProgram’, ‘-searchTime’, ‘-searchDepth’,
‘-saveGameFile’, ‘-autosave’, ‘-savePositionFile’ and ‘-boardSize’. See Chap-

ter 3 [Invocation], page 5.

7.2 Starting a CMail game.

Type ‘cmail’ from a shell to start a game as white. After an opening message, you will be prompted for a game name, which is optional – if you simply press return, the game name will take the form ‘you-VS-opponent’. You will next be prompted for the short name of your opponent. If you haven’t played this person before, you will also be prompted for his/her email address. ‘cmail’ will then invoke XBoard in the background. Make your first move and select ‘Mail Move’ from the ‘File’ menu. See Section 4.1 [File Menu], page 16. If all is well, ‘cmail’ will mail a copy of the move to your opponent. If you select ‘Exit’ without having selected ‘Mail Move’ then no move will be made.

7.3 Answering a move.

When you receive a message from an opponent containing a move in one of your games, simply pipe the message through ‘cmail’. In some mailers this is as simple as typing | cmail when viewing the message, while in others you may have to save the message to a file and do cmail < file at the command line. In either case ‘cmail’ will display the game using XBoard. If you didn’t exit XBoard when you made your first move then ‘cmail’ will do its best to use the existing XBoard instead of starting a new one. As before, simply make a move and select ‘Mail Move’ from the ‘File’ menu. See Section 4.1 [File Menu], page 16. ‘cmail’ will try to use the XBoard that was most recently used to display the current game. This means that many games can be in progress simultaneously, each with its own active XBoard.

If you want to look at the history or explore a variation, go ahead, but you must return to the current position before XBoard will allow you to mail a move. If you edit the game’s history you must select ‘Reload Same Game’ from the ‘File’ menu to get back to the original position, then make the move you want and select ‘Mail Move’. As before, if you decide you aren’t ready to make a move just yet you can either select ‘Exit’ without sending a move or just leave XBoard running until you are ready.

Because XBoard can now detect checkmate and stalemate, ‘cmail’ now handles game termination sensibly. As well as resignation, the ‘Action’ menu now allows draws to be offered and accepted for ‘cmail’ games.

For multi-game messages, only unfinished and just-finished games will be included in email messages. When all the games are finished, they are archived in the user’s archive directory, and similarly in the opponent’s when he or she pipes the final message through ‘cmail’. The archive file name includes the date the game was started.

It’s possible to have a ‘cmail’ message carry more than one game. This feature was implemented to handle IECG (International Email Chess Group) matches, where a match consists of 1 game as white and 1 as black, with moves transmitted simultaneously. In case there are more general uses, ‘cmail’ itself places no limit on the number of black/white games contained in a message; however, XBoard does.

7.4 Known CMail problems.

It’s possible that a strange conjunction of conditions may occasionally mean that ‘cmail’ has trouble reactivating an existing XBoard. If this should happen, simply trying it again should work.

If not, remove the file that stores the XBoard’s PID (‘game.pid’) or use the ‘-xreuse’ option to force ‘cmail’ to start a new XBoard.

Versions of ‘cmail’ after 2.16 no longer understand the old file format that XBoard used to use and so cannot be used to correspond with anyone using an older version.

Versions of ‘cmail’ older than 2.11 do not handle multi-game messages, so multi-game correspondence is not possible with opponents using an older version.

8 Environment variables

Game and position files are found in a directory named by the CHESSDIR environment variable. If this variable is not set, the current working directory is used. If CHESSDIR is set, XBoard actually changes its working directory to $CHESSDIR, so GNU Chess listing files will also be stored there as well.

9 Known limitations and bugs

There is no way for two people running copies of XBoard to play each other without going

through the Internet Chess Server.

Under some circumstances, your ICS password may be echoed when you log on.

If you are connecting to the ICS by running telnet on an Internet provider or firewall host, you may find that each line you type is echoed back an extra time after you hit Return. If your Internet provider is a Unix system, you can probably turn its echo off by typing

     stty -echo

after you log in, and/or typing ^E-RETURN (CONTROL-E followed by the RETURN key) to the telnet program after you have logged into ICS. It is a good idea to do this if you can, because the extra echo can occasionally confuse XBoard’s parsing routines.

The game parser recognizes only algebraic notation.

The internal move legality tester does not look at the game history, so in some cases it misses illegal castling or en passant captures. It permits castling with the king on the d file because this is possible in some "wild 1" games on ICS. It does not check piece drops in bughouse to see if you actually hold the piece you are trying to drop. However, if you attempt an illegal move when using GNU Chess (or the ICS), XBoard will accept the error message that comes back, undo the move, and let you try another.

FEN positions saved by XBoard do not include correct information about whether castling or en passant are legal.

The mate detector does not understand that non-contact mate is not really mate in bughouse. The only problem this causes while playing is minor: a ‘#’ (mate indicator) character will show up after a non-contact mating move in the move list. XBoard will not assume the game is over at that point.

‘Edit Game’ mode always uses non-bughouse rules. Although you can load and edit games that contain piece drops, the piece menus are not active, so you cannot insert piece drops. Also, XBoard thinks an edited game is over when a mating move is inserted, even if the mate is non-contact.

Some XBoard functions may not work with versions of GNU Chess earlier than 4.0, patchlevel 73, or with versions of Crafty other than the "xboard-supported" version; see GETTING CRAFTY above. A few functions work with GNU Chess but not Crafty, or vice versa.

The menus may not work if your keyboard is in Caps Lock mode. This seems to be a problem with the Athena menu widget.

XBoard depends on the Xt Intrinsics and Athena Widget Set of X11R4 or later. In particular, X11R3 just won’t do.

10 Reporting problems

Report bugs and problems with XBoard to

bug-gnu-chess@prep.ai.mit.edu.

Please use the ‘script’ program to start a typescript, run XBoard with the ‘-debug’ option, and include the typescript output in your message. Also tell us what kind of machine and what operating system version you are using. The command ‘uname -a’ will often tell you this. Here is a sample of approximately what you should type:

script
uname -a
./configure
make
./xboard -debug
exit
mail bug-gnu-chess@prep.ai.mit.edu
Subject: problem with foobaz command in XBoard
Description of the problem
~r typescript
.

If you improve XBoard, please send a message about your changes to bug-gnu-chess, and we will get in touch with you about merging them in to the main line of development.

Send CMail bug reports/suggestions to

evan@quadstone.co.uk

11 Authors and contributors to XBoard

Tim Mann (mann@pa.dec.com)

has been responsible for XBoard versions 1.3 through 3.5 and for WinBoard, a port of XBoard to Microsoft Win32 (Windows NT and Windows 95)

Jochen Wiedmann (wiedmann@neckar-alb.de)

ported XBoard to the Amiga, creating AmyBoard, and converted the documentation to texinfo. He was responsible for AmyBoard versions through 330.5 (based on xboard 3.3.pl0).

Frank McIngvale (frankm@hiwaay.net)

added "Click-Click" mode, the Analysis modes, piece flashing, ZIICS import, and ICS text colorization.

Chris Sears and Dan Sears wrote the original XBoard. They were responsible for versions 1.0 through 1.2.

Elmar Bartel contributed the new piece bitmaps introduced in version 3.2.

Evan Welsh (evan@quadstone.co.uk)

wrote CMail.

Patrick Surry helped with design, testing and documenting CMail.

John Chanak contributed the initial implementation of ICS mode.

Wayne Christopher created XChess; the color scheme and the old 80x80 piece bitmaps were taken from it.

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you changed the

files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part contains

or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it. Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program.

In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code, which must

be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WAR-RANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFOR-MANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DE-FECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CON-SEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSI-BILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it starts in an interactive

mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright
interest in the program `Gnomovision'
(which makes passes at compilers) written
by James Hacker.

signature of Ty Coon, 1 April 1989

Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking GNU GENERAL PUBLIC LICENSE 43 proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Index 44

Index

.

.cmailaliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 .cmailgames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 .icsrc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

>

>, Button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 >>, Button. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

<

<, Button . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 <<, Button. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

A

abort, ICS command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Abort, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 About XBoard, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 23 accept, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Accept, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Action, Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 adjourn, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Adjourn, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Always Queen, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . 21 alwaysPromoteToQueen, option . . . . . . . . . . . . . . . . . . . . 11 Analysis Mode, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 17 Analyze File, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Auto Comment, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 21 Auto Flag, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Auto Observe, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Auto Save, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 autobs, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autoCallFlag, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autocomm, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autoComment, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autoflag, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autoObserve, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 autosave, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 autoSaveGames, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

B

Back to Start, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Backward, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Bell, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 bell, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 bitmapDirectory, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 blackPieceColor, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 blackString, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 bm, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 board size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 boardSize, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Book, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 borderXoffset, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 borderYoffset, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Bug reports. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34, 35

C

Call Flag, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 checkLegality, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Chess program options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 CHESSDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 clickClick, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 clock, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 clockFont, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 clockMode, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 cmail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 colorChallenge, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorChannel, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorChannel1, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorize, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorKibitz, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorNormal, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorShout, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorSShout, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 colorTell, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Contributors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 coordFont, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 coords, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

D

darkSquareColor, option . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 debug, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 debugMode, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 decline, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Decline, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 display, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 draw, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Draw, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

E

Edit Comment, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 19 Edit Game, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Edit Position, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Edit Tags, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 examine, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Exit, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

F

fcp, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 fh, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 File Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 finger, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 firstChessProgram, option. . . . . . . . . . . . . . . . . . . . . . . . . . . 6 firstHost, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 flashCount, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 flashRate, rate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Flip View, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 flip, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 flipView, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Font, clock . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Font, coordinates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 font, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Forward to End, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . 21 Forward, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

G

games, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 gateway, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 geometry, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Get Move List, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . 22 getMoveList, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

H

Help Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 help, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Hint, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

I

I/O options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 iconic, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ICS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ICS Client, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ICS Input Box, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . 19 ICS options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ICS, addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ics, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 icscomm, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 icshost, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 icsinput, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 icslogon, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 icsport, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 inc, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Info XBoard, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 initString, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Internet Chess Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 internetChessServerCommPort, option . . . . . . . . . . . . . . 9 internetChessServerHost, option . . . . . . . . . . . . . . . . . . . . . 9 internetChessServerInputBox, option . . . . . . . . . . . . . . . . 8 internetChessServerLogonScript, option . . . . . . . . . . . . . 7 internetChessServerMode, option . . . . . . . . . . . . . . . . . . . . 7 internetChessServerPort, option . . . . . . . . . . . . . . . . . . . . . 9 introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

K

Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

L

legal, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 lgf, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 lgi, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 lightSquareColor, option . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Load Game, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Load Next Game, Menuitem . . . . . . . . . . . . . . . . . . . . . . . 16 Load Position, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Load Previous Game, Menuitem . . . . . . . . . . . . . . . . . . . 16 loadGameFile, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 loadGameIndex, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 loadPositionFile, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 loadPositionIndex, option . . . . . . . . . . . . . . . . . . . . . . . . . . 11 lpf, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 lpi, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

M

Machine Black, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . 17 Machine White, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 17 Mail Move, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Man XBoard, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 matchMode, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Menu, Action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Menu, File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Menu, Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Menu, Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Menu, Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Menu, Step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 mm, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Mode Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 mono, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 monoMode, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Move Now, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 moves, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 movesPerSession, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 mps, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 msLoginDelay, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

N

ncp, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 noChessProgram, option . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

O

observe, ICS command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Old Save Style, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . 22 oldmoves, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 oldsave, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 oldSaveStyle, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Options Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 options, Chess program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Options, I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Options, ICS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Options, miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Options, User interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

P

Pause, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Periodic Updates, Menuitem . . . . . . . . . . . . . . . . . . . . . . . 22 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Q

queen, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Quiet Play, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 quiet, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 quietPlay, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

R

Reload CMail Message, Menuitem. . . . . . . . . . . . . . . . . . 17 Reload Same Game, Menuitem . . . . . . . . . . . . . . . . . . . . . 16 remoteShell, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 remoteUser, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Reporting problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Reset, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Resign, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Retract Move, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Revert, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 ringBellAfterMoves, option. . . . . . . . . . . . . . . . . . . . . . . . . 11 rsh, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ruser, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

S

Save Game, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 saveGameFile, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 savePositionFile, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 scp, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 sd, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 searchDepth, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 searchTime, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 secondChessProgram, option . . . . . . . . . . . . . . . . . . . . . . . . 6 secondHost, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 set, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 sgf, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 sh, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Shortcut keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Show Coords, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Show Game List, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . 18 Show Thinking, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 22 showCoords, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 showThinking, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 size, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 spf, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 st, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Step Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Stop Examining, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . 20 Stop Observing, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . 20

T

tc, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 td, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 telnet, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 telnetProgram, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 thinking, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 timeControl, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 timeDelay, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 timeIncrement, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 title, option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 titleInWindow, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Truncate Game, Menuitem. . . . . . . . . . . . . . . . . . . . . . . . . 21 Two Machines, Menuitem . . . . . . . . . . . . . . . . . . . . . . . . . . 17

U

User interface options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 useTelnet, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

V

vars, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

W

whitePieceColor, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 whiteString, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 who, ICS command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

X

xpm, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Other programs you can use with XBoard. . . . . . . . . 2

1.1 GNU Chess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Crafty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 zic2xpm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 How to configure, compile and install XBoard …. 4

3 Command options supported by XBoard . . . . . . . . . . 5

3.1 Controlling the chess program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Connecting to and using ICS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Load and Save options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.4 Look and Feel options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.5 Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

4 Menus, buttons and keys . . . . . . . . . . . . . . . . . . . . . . . . . 16

4.1 Accessing external games and positions . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.2 Selecting XBoard’s mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.3 Talking to GNU chess or ICS opponents . . . . . . . . . . . . . . . . . . . . . . . . 19 4.4 Controlling the game . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.5 User Preferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.6 Getting help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.7 Other shortcut keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

5 Using XBoard with an Internet Chess Server (ICS) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

6 Connecting to the ICS through a firewall . . . . . . . . 27

7 Using XBoard for electronic correspondence chess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

7.1 Invoking CMail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 7.2 Starting a CMail game. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.3 Answering a move. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.4 Known CMail problems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

8 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

9 Known limitations and bugs . . . . . . . . . . . . . . . . . . . . . . 34 10 Reporting problems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

11 Authors and contributors to XBoard. . . . . . . . . . . . 36

GNU GENERAL PUBLIC LICENSE. . . . . . . . . . . . . . . . 37

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND

MODIFICATION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 How to Apply These Terms to Your New Programs . . . . . . . . . . . . . . . . . . . 42

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44